# Sky-Drive: A Distributed Multi-Agent Simulation Platform for Socially-Aware and Human-AI Collaborative Future Transportation

### v1

@@ V-B HAIM-based Deep Reinforcement Learning
TABLE II: The performance of PPO, HACO, and HAIM-DRL methods.
Method Test Safety Violation Test Return Test Disturbance Rate Test Success Rate Train Samples
PPO 80.84 1591.00 - 0.35 500,000
HACO 12.14 1578.43 0.0137 0.35 8,000
HAIM-DRL 11.25 1590.85 0.0121 0.38 8,000
@@
TABLE III: Performance comparison of PE-RLHF with different physics-based model combinations.
Method Driving Operation Training Testing
Total Safety Violation $\downarrow$ Episodic Return $\uparrow$ Success Rate (%) $\uparrow$ Safety Violation $\downarrow$ Travel Distance $\uparrow$ Travel Velocity $\uparrow$ Total Overtake Count $\uparrow$
IDM-MOBIL Longitudinal & Lateral - 206.30 $\pm$35.23 0.31 $\pm$0.15 0.49 $\pm$0.08 108.56 $\pm$55.23 19.78 $\pm$2.67 0 $\pm$0
PE-RLHF (without) - 39.45 $\pm$ 12.32 302.67 $\pm$ 21.88 0.73 $\pm$ 0.05 1.48 $\pm$ 0.43 138.23 $\pm$ 4.28 16.58 $\pm$ 0.96 6.14 $\pm$ 1.12
PE-RLHF (with IDM) Longitudinal only 28.79 $\pm$ 9.97 348.52 $\pm$ 19.67 0.79 $\pm$ 0.03 0.98 $\pm$ 0.29 149.87 $\pm$ 4.10 18.92 $\pm$ 0.94 7.83 $\pm$ 1.03
PE-RLHF (with MOBIL) Lateral only 21.56 $\pm$ 8.54 368.11 $\pm$ 18.45 0.81 $\pm$ 0.04 0.74 $\pm$ 0.19 159.34 $\pm$ 3.14 20.43 $\pm$ 0.51 9.76 $\pm$ 1.17
PE-RLHF (with IDM-MOBIL) Longitudinal & Lateral 16.61 $\pm$ 9.96 391.48 $\pm$ 20.47 0.85 $\pm$ 0.04 0.47 $\pm$ 0.01 177.00 $\pm$ 3.74 21.85 $\pm$ 0.02 16.33 $\pm$ 4.61
@@
V-C Physics-enhanced Reinforcement Learning with Human Feedback
To validate the integration of knowledge in transportation science within Sky-Drive, we implemented and tested the physics-enhanced reinforcement learning with human feedback (PE-RLHF) [57]. Unlike traditional methods that may falter with imperfect human feedback, PE-RLHF establishes a trustworthy safety performance lower bound through well-established traffic flow models.
PE-RLHF implements this idea through three policies: a human policy $\pi_{\text{human}}$, a physics-based policy $\pi_{\text{phy}}$ derived from traffic flow models, and a learning agent policy $\pi_{\text{AV}}$. When no takeover occurs, the AV agent executes $\pi_{\text{AV}}$ and updates its policy through environment-driven exploration. When human intervention is detected, the PE-RLHF compares the expected values of actions suggested by the human and the physics-based policy.
$\mathcal{T}_{\text{PE-HAI}}(s)=\begin{cases}a_{\text{hybrid}},&\text{if takeover}\\ a_{\text{AV}},&\text{otherwise}\end{cases}$ (4)
In detail, if the human action is expected to yield better outcomes, it is selected; otherwise, the physics-based action is executed. This hybrid policy guarantees that decisions adhere to a minimum standard of safety and efficiency:
$a_{\text{hybrid}}$ $=T_{\text{select}}(s)$ (5)
=$=\begin{cases}a_{\text{human}},&\text{if }Q^{\phi}(s,a)-Q^{\phi}(s,a)\geq\varepsilon_{\text{select}}\\ a_{\text{phy}},&\text{otherwise}\end{cases}$
The Eq. (5) ensures that the system always executes the action with higher expected value, establishing a performance floor guaranteed by interpretable physics-based models, even when human feedback quality deteriorates.
The overall learning objective of PE-RLHF is formulated as [57]:
$\max_{\pi}\mathbb{E}\left[\psi\hat{Q}(s_{t},a_{t}^{\text{hybrid}})-\alpha\log\pi_{\text{AV}}(a_{t}^{\text{AV}}|s_{t};\theta)-\beta Q^{\text{int}}(s_{t},a_{t}^{\text{AV}})\right]$ (6)
where $\hat{Q}(s_{t},a_{t}^{\text{hybrid}})$ is a proxy value function, the entropy term encourages exploration, and $Q^{\text{int}}(s_{t},a_{t}^{\text{AV}})$ minimizes the need for human intervention.
Tab. III presents the performance comparison of PE-RLHF under different configurations, including standalone physics-based control (IDM-MOBIL) and PE-RLHF variants with or without integrated physics models. The full PE-RLHF configuration consistently outperforms all baselines across all stages. PE-RLHF improves episodic return and success rate by nearly 90% compared to the standalone model. Meanwhile, it achieves the lowest safety violation (0.47) and the longest travel distance (177.00 m). Moreover, the full configuration reaches the highest travel velocity (21.85 km/h) and completes the most overtaking maneuvers (16.33), whereas the IDM-MOBIL baseline fails to overtake at all. These results validate Sky-Drive’s ability to support the integration of physics knowledge and human feedback for learning safe and efficient autonomous driving policies.
V-D Vision Language Model-Enabled Reinforcement Learning
TABLE IV: Performance comparison with baselines during testing. Mean and standard deviation over 3 seeds.
VLM-SR 0.53 $\pm$ 0.27 0.02 $\pm$ 0.00 47.9 $\pm$ 9.2 0.18 $\pm$ 0.25 0.0 $\pm$ 0.0
LORD 0.17 $\pm$ 0.08 0.02 $\pm$ 0.02 45.1 $\pm$ 57.1 0.02 $\pm$ 0.02 0.0 $\pm$ 0.0
@@ V-D Vision Language Model-Enabled Reinforcement Learning
As shown in Tab. IV, VLM-RL significantly outperforms existing approaches across all key metrics. VLM-RL achieves the highest success rate and route completion, while maintaining a low collision speed of 0.02 km/h—matching the safety level of the most conservative baselines. Unlike existing VLM-based methods, which suffer from overly cautious behavior and near-zero task success, VLM-RL balances safety with efficiency, reaching an average speed of 19.3 km/h and a total driving distance of 2028.2 m. Compared to strong LLM-based methods such as Revolve, VLM-RL maintains comparable success and completion rates while drastically reducing collision speed. The successful implementation of VLM-RL within the Sky-Drive platform validates its capability to support large-scale, multimodal policy learning grounded in human-understandable semantics.
V-E Personalized Safety-Critical Curriculum Learning
TABLE V: Performance comparison with baselines in the safety-critical test scenarios.
@@ V-E Personalized Safety-Critical Curriculum Learning
As shown in Tab. V, CurricuVLM achieves the best overall performance across all key metrics, demonstrating both high safety and training effectiveness. In terms of task performance, CurricuVLM achieves the highest episode reward (48.9) and road completion rate (73.4%), while maintaining a low crash rate (25.1%), outperforming strong baselines such as CAT and CLIC. It also records the highest total driving distance (48.4m) and failure-to-success rate (39.1%), indicating superior adaptability to previously failed scenarios. Meanwhile, its success-to-success rate (73.5%) reflects strong behavioral consistency and learning stability. These results validate that CurricuVLM not only enhances policy robustness under long-tail safety-critical scenarios, but also integrates seamlessly into Sky-Drive’s human-AI collaboration mechanism.
V-F Interactive and Editable Traffic Scenario Generation
@@
To validate Sky-Drive’s capability in supporting multimodal, human-centered scenario generation, we implemented the Talk2Traffic [61], which enables intuitive and editable traffic scenario creation through natural language, speech, and sketch-based inputs, as illustrated in Fig. 5. It demonstrates Sky-Drive’s support for seamless integration of multimodal large language models (MLLMs) into simulation workflows for autonomous driving research.
The core idea of Talk2Traffic is to bridge the gap between human designers’ intent and executable traffic simulation. The system first interprets user instructions through a multimodal encoder [61]:
$\mathbf{z}=\text{MLLM}(p,l,s),$ (11)
where $p$ is the task prompt, $l$ represents textual or spoken language input, and $s$ denotes sketch-based spatial constraints. The output $\mathbf{z}$ is a structured scene representation capturing key elements such as map layout, agent behaviors, and environmental conditions.
To generate executable code, Talk2Traffic adopts a retrieval-augmented generation (RAG) strategy based on a curated database of description-code pairs:
$\mathcal{D}=\{(d_{j},c_{j})|j\in\{1,\ldots,m\}\},$ (12)
where $d_{j}$ is a natural language description and $c_{j}$ is the corresponding Scenic code snippet. Relevant snippets are retrieved using semantic similarity. A key feature of Talk2Traffic is its interactive refinement mechanism, where users can iteratively modify the scenario based on feedback.
Experimental results demonstrate Talk2Traffic’s superior capability in generating diverse and challenging scenarios. It achieves the lowest average collision rate (0.877) across all scenario types, surpassing the next-best baseline by 4.6%. In high-complexity interactions such as red light running (0.900) and unprotected left turn (0.833), Talk2Traffic consistently outperforms other methods. Some qualitative examples of generated scenarios are illustrated in Fig. 5. By integrating Talk2Traffic into the AIHM module, Sky-Drive enables dynamic and goal-directed scenario generation, allowing personalized training and evaluation of autonomous agents under richly varied, realistic, and safety-critical conditions.
V-G Accident Data Replay
@@ VI-A3 Traffic Brains
Sky-Drive will position foundation models as intelligent “traffic brains” that govern decision-making in complex, multi-agent traffic environments. While general-purpose models such as Qwen [23], GPT-4 [21], and Llama [22] exhibit strong language and reasoning abilities, they will require domain-specific adaptation to meet the demands of autonomous driving. To address this challenge, Sky-Drive will leverage transportation-specific datasets—including LMDrive [67], CCD [68], DoTA [69], and DriveCoT [70]—to fine-tune pre-trained foundation models. This fine-tuning pipeline is designed to enhance the model’s ability to handle dynamic scenario adaptation, hierarchical reasoning, and multitask decision-making, including generating safe control actions (e.g., steering, throttle, and braking) and predicting critical safety metrics such as time-to-collision (TTC). The refined models will be deployed within the Sky-Drive simulation environment to enable coordinated behavior across AVs and other components, facilitating holistic control and system-level optimization.
@@ VI-B1 Simulation-to-Reality Integration
As shown in Fig. 1 (e), the center of the HIL framework is a Ford E-Transit electric van retrofitted with fully automated driving capabilities. The vehicle is equipped with a comprehensive sensor suite—including LiDAR, radar, high-resolution cameras, and OxTS navigation units—and operates on a drive-by-wire system connected to an industrial-grade computing rack. The vehicle’s software stack builds upon ROS-based open-source packages, further augmented with proprietary Sky-Drive modules to support advanced planning, control, and communication tasks. To complement the vehicle-level testing, Sky-Drive also deploys portable roadside infrastructure units outfitted with traffic lights, regulatory signage, cameras, and LiDAR systems. These roadside units enable systematic evaluation of V2X communication and cooperative perception algorithms across diverse environmental settings [71, 72]. By maintaining connectivity with Sky-Drive’s DT environment, the framework allows algorithms to be first validated in simulation and then deployed to physical vehicles with minimal transition cost, significantly accelerating development cycles while ensuring safe real-world performance.
VI-B2 Teleoperated Driving
The HIL framework also establishes a solid foundation for developing and testing teleoperated driving. Teleoperated driving allows humans (teleoperators) to remotely control vehicles, particularly in challenging scenarios, complementing fully/highly autonomous solutions. It is one of the important use cases of V2X communication, specified in the 3GPP standards [73]. Sky-Drive’s ROS integration enables wireless connectivity between its testbed vehicle and human-in-the-loop simulation platform—operated by a teleoperator—via cellular or satellite networks, e.g., 5G. Considering the wild fluctuations of network bandwidth, round-trip time (RTT), jitter time, and packet loss under driving conditions of 5G [74], Sky-Drive facilitates the collaboration between the vehicle and the simulation platform to dynamically decide what data (RGB images, LiDAR point cloud, and/or their pre-processed data) to transmit and how to transmit them to meet the end-to-end latency requirement for the teleoperation, i.e., below 100 milliseconds [75].

### v2

@@ V-B HAIM-based Deep Reinforcement Learning
TABLE IV: Performance comparison with baselines in the safety-critical test scenarios.
@@
V-C Vision Language Model-Enabled Reinforcement Learning
@@
As shown in Tab. III, VLM-RL significantly outperforms existing approaches across all key metrics. VLM-RL achieves the highest success rate and route completion, while maintaining a low collision speed of 0.02 km/h—matching the safety level of the most conservative baselines. Unlike existing VLM-based methods, which suffer from overly cautious behavior and near-zero task success, VLM-RL balances safety with efficiency, reaching an average speed of 19.3 km/h and a total driving distance of 2028.2 m. Compared to LLM-based methods such as Revolve, VLM-RL maintains comparable success and completion rates while drastically reducing collision speed. The successful implementation of VLM-RL within the Sky-Drive platform validates its capability to support large-scale, multimodal policy learning grounded in human-understandable semantics.
V-D Personalized Safety-Critical Curriculum Learning
@@
As shown in Tab. IV, CurricuVLM achieves the best overall performance across all key metrics, demonstrating both high safety and training effectiveness. In terms of task performance, CurricuVLM achieves the highest episode reward (48.9) and road completion rate (73.4%), while maintaining a low crash rate (25.1%), outperforming baselines such as CAT and CLIC. It also records the highest total driving distance (48.4m) and failure-to-success rate (39.1%), indicating superior adaptability to previously failed scenarios. Meanwhile, its success-to-success rate (73.5%) reflects strong behavioral consistency and learning stability. These results validate that CurricuVLM not only enhances policy robustness under long-tail safety-critical scenarios, but also integrates seamlessly into Sky-Drive’s human-AI collaboration mechanism. Some qualitative examples of generated scenarios are illustrated in Fig. 5.
@@
V-E Accident Data Replay
@@ VI-A3 Traffic Brains
Sky-Drive will position foundation models as intelligent “traffic brains” that govern decision-making in complex, multi-agent traffic environments. While general-purpose models such as Qwen [24], GPT [22], and Llama [23] exhibit strong language and reasoning abilities, they will require domain-specific adaptation to meet the demands of autonomous driving. To address this challenge, Sky-Drive will leverage transportation-specific datasets—including LMDrive [68], CCD [69], DoTA [70], and DriveCoT [71]—to fine-tune pre-trained foundation models. This fine-tuning pipeline is designed to enhance the model’s ability to handle dynamic scenario adaptation, hierarchical reasoning, and multitask decision-making, including generating safe control actions (e.g., steering, throttle, and braking) and predicting critical safety metrics such as time-to-collision (TTC). The refined models will be deployed within the Sky-Drive simulation environment to enable coordinated behavior across AVs and other components, facilitating holistic control and system-level optimization.
@@ VI-B1 Simulation-to-Reality Integration
As shown in Fig. 1 (e), the center of the HIL framework is a Ford E-Transit electric van retrofitted with fully automated driving capabilities. The vehicle is equipped with a comprehensive sensor suite—including LiDAR, radar, high-resolution cameras, and OxTS navigation units—and operates on a drive-by-wire system connected to an industrial-grade computing rack. Sky-Drive connects its real-time simulation environment with the physical vehicle through standardized ROS interfaces to realize hardware-in-the-loop testing. Simulated sensor streams (e.g., camera and LiDAR) are fed into the vehicle’s onboard systems, which process the data and return control commands (e.g., throttle, brake, steering) to the simulator. This closed-loop integration allows developers to evaluate real hardware behavior under diverse and repeatable traffic scenarios, including safety-critical scenarios that are difficult or unsafe to test on public roads. By testing algorithms in simulation before deployment, Sky-Drive can bridge the sim-to-real gap and ensure that experimental validation can be conducted safely without exposing operators or the public to real-world risks.
VI-B2 Remote Driving and Data Collection
The HIL framework also establishes a solid foundation for developing and testing teleoperated driving. Teleoperated driving allows humans (teleoperators) to remotely control vehicles, particularly in challenging scenarios, complementing fully/highly autonomous solutions. It is one of the important use cases of V2X communication, specified in the 3GPP standards [72, 73, 74]. Sky-Drive’s ROS integration enables wireless connectivity between the testbed vehicle and the human-in-the-loop framework via cellular or satellite networks, such as 5G. To support this, Sky-Drive integrates a 5G mmWave mobile hotspot (i.e., NETGEAR Nighthawk M6 Pro) to enable remote control and real-time data collection. This setup allows researchers to control vehicles without being physically onboard, ensuring operator safety during high-risk scenarios. Meanwhile, it supports remote collection of human behavioral data (e.g., steering patterns and pedal inputs), even when the vehicle and experiment coordinators are located in different cities or campuses.
@@ VII Conclusions
Acknowledgments
This work was funded by the U.S. Department of Transportation, Award #69A3552348305. In addition, this research was supported by grants from NVIDIA and utilized NVIDIA RTX PRO 6000 Blackwell Max-Q Workstation Edition GPUs and A100 GPU-Hours on Saturn Cloud.